Three identical coils, each of resistance 15 Ω and inductance 42 mH are
connected (a) in star and (b) in delta to a 415 V, 50 Hz, 3-phase supply.
Determine the total power dissipated in each case.
1
Expert's answer
2020-12-06T17:18:58-0500

(a) Star connection

Inductive reactance

"X_L = 2πfL \\

X_L = 2π \times 50 \times 42 \times 10^{-3} = 13.19 \;Ω"

Phase impedance

"Z_{ph} = \sqrt{R^2 + X_L^2} \\

Z_{ph} = \sqrt{15^2 + 13.19^2} = 19.97"

Line voltage

"V_{line} = \sqrt{3}V_{ph} \\

V_{ph} = \frac{V_{line}}{\sqrt{3}} = \frac{415}{\sqrt{3}} = 240 \;volts"

Phase current

"I_{ph } = \frac{V_{ph}}{Z_{ph}} \\

I_{ph} = \frac{240}{19.97} = 12.02 \;A"

Line current

"I_{line} = I_{ph} = 12.02 \;A"

Dissipated power

"P = 3I_{line}^2R \\

P = 3 \times (12.02)^2 \times 15 = 6.6 \;kW"

(b) Delta connection

"V_{line} = V_{ph} = 415 \;V \\

Z_{ph} = 19.97 \\

I_{ph} = \frac{V_{ph}}{Z_{ph}} \\

I_{ph} = \frac{415}{19.97} = 20.78 \;A"

Dissipated power

"P = \sqrt{3}I_{ph}^2R \\

P = \sqrt{3} \times 20.78^2 \times 15 = 11.22 \;kW"